OpenStudy (anonymous):

Plz Help will give Medal and Fan.. The two numbers rolled can be added to get a sum. Find P(sum is even).

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

Choices are 1/4 15/36 1/2 3/4

OpenStudy (anonymous):

can you help? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you'll use that table again in each box in the table, I want you to add the coordinates of each ordered pair for instance, (1,2) will have 3 in that box because 1+2 = 3

jimthompson5910 (jim_thompson5910):

fill out the entire table like this

OpenStudy (anonymous):

so I just add everything like 4,4 = 4+4=8

jimthompson5910 (jim_thompson5910):

yes, for each ordered pair to see what the sum of the dice is

OpenStudy (anonymous):

ok got it

jimthompson5910 (jim_thompson5910):

tell me when you've done that

OpenStudy (anonymous):

Here u go.. @jim_thompson5910

OpenStudy (anonymous):

SO I need to list the even numbers

jimthompson5910 (jim_thompson5910):

yes, you need to count how many times you get an even sum

OpenStudy (anonymous):

Yes ok

OpenStudy (anonymous):

18 even numbers

jimthompson5910 (jim_thompson5910):

out of 36 total, so 18/36 = 1/2

OpenStudy (anonymous):

yeah I was just about to say that

OpenStudy (anonymous):

Can you help me with other questions more

jimthompson5910 (jim_thompson5910):

I'll help with one more